### v1

@@ 3 Disscusion
The development of XiChen represents a significant breakthrough in fully AI-driven global weather forecasting. Leveraging the scalability of foundation models, XiChen is constructed by a pretrained medium-range forecasting model and fine-tuned to serve as observation operators and DA models. The entire pipeline of XiChen, from DA to 10-day weather forecasting, can be accomplished within only 17 seconds, representing an exceeding 400-fold acceleration compared to the operational NWP system [17]. By integrating GDAS prepbufr, AMSU-A, and MHS observations, XiChen offers data-driven global weather forecasting capabilities and demonstrates superior performance compared to the GFS in both DA and forecasting tasks for several key variables. Notably, XiChen achieves a skillful lead time of over 8.25 days, which is comparable to that of the GFS while surpassing Aardvark [22], and GraphDOP [23]. Moreover, XiChen exhibits high scalability: when new satellite observations are introduced, it requires only fine-tuning the foundation model to the respective observation operator, followed by fine-tuning the corresponding assimilation model using the cascade DA framework proposed in this study. Similarly, if an observation becomes temporarily unavailable, the assimilation process can exclude the corresponding model from the pipeline. These capabilities are critical for advancing AI-based DA and forecasting toward practical operationalization. This achievement underscores the transformative potential of AI in advancing operational weather forecasting and highlights the necessity of sustained investments in AI research to tackle the complexities of Earth system modeling. From a multi-agent perspective [41], the XiChen system can be conceptualized as a weak multi-agent system leveraging the pre-training and fine-tuning paradigms of Foundation Models. Within this framework, multiple specialized “agents” are deployed to address distinct tasks: predicting future atmospheric variations based on current states, mapping current atmospheric conditions to satellite observation spaces, and assimilating observations to correct forecast states. Each of these “agents” operates autonomously within its designated function, collectively forming a fully AI-driven global forecasting system that spans the core components of the NWP pipeline—-from assimilating diverse observations to global weather forecasts. As weather forecasting enters a new era, XiChen establishes a robust foundation for applying AI to deepen our forecasting of planetary systems, ultimately guiding decisions with profound implications for society and the environment. However, this does not imply that AI-driven global weather forecasting systems can fully replace NWP. Most AI-driven weather forecasting systems, including XiChen, rely heavily on reanalysis fields produced by NWP systems for training. Consequently, fully AI-driven global weather forecasting, exemplified by XiChen, serves as a valuable complement to NWP rather than a replacement.
@@
Although XiChen demonstrates significant potential for practical applications, it also exhibits certain limitations. First, the model experiences smoothing effects during both the DA and weather forecasting processes, a common challenge in AI-based weather forecasting models [42]. Addressing this issue requires further dedicated research. Potential solutions include designing more suitable loss functions [43], incorporating physical constraints [44], or employing post-processing techniques based on generative models [40]. Second, the current version of XiChen has been trained and tested exclusively at a resolution of 1.40625. To predict finer-scale weather phenomena, future work should focus on extending this resolution to higher levels, such as 0.25 [9] or even 0.09 [45]. Third, the current implementation of XiChen integrates indirect observations from only two satellite instruments, AMSU-A and MHS, which is substantially fewer than the number of satellites used in operational NWP systems [40]. In the future, incorporating additional indirect observational data could significantly enhance the performance of the analysis field. Fourth, when MHS observations are assimilated following the GDAS prepbufr, slight increases in upper-atmosphere temperature and geopotential errors are evident. These discrepancies primarily arise because MHS brightness temperature observations are affected by clouds, precipitation, and surface radiation—factors that XiChen does not currently account for. Additionally, the substantial adjustments to specific humidity during the DA process may induce imbalances. In the future, addressing this issue may involve incorporating additional relevant variables and embedding physical equilibrium constraints into XiChen as potential solutions. Finally, the model produces only deterministic DA and weather forecasts. However, making critical decisions requires awareness of all possible weather scenarios and their associated probabilities. Therefore, it is necessary to extend the model to a probabilistic version. Addressing this limitation in the future may involve reconstructing the model to adopt a probabilistic framework [46] or incorporating perturbations through the deterministic XiChen system [47]. Ultimately, our goal is to develop a fully AI-driven global weather forecasting system that rivals the IFS HRES model in both assimilation and forecasting capabilities. Achieving this ambitious goal necessitates expertise in atmospheric science and the seamless integration of strengths from the atmospheric and AI communities to work together.
@@ 4.1 Notation and problem statement
We aim to develop a fully AI-driven global weather forecasting system that integrates both forecasting and DA models. For the forecasting model, given a system state $\mathbf{x}(t)$, our objective is to predict the state $\mathbf{x}(t+\Delta t)$ at a future time $t+\Delta t$. To accomplish this, the forecasting model $\mathcal{M}_{t\rightarrow t+\Delta t}:\mathbf{x}(t)\mapsto\hat{\mathbf{x}}(t+\Delta t)$ is trained to map the state of the atmosphere at present to a predicted state $\hat{\mathbf{x}}(t+\Delta t)$ for the future time $t+\Delta t$. Regarding the assimilation model, given a background field $\mathbf{x}^{b}(t_{0})$ at time $t_{0}$, GDAS prepbufr observations $\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\}$, and satellite observations $\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}$ during the DAW $[t_{0},\cdots,t_{K}]$, our goal is to learn to estimate the optimal initial field $\mathbf{x}^{a}(t_{0})$ by fusing these data. To assimilate the satellite observations, we need to learn an observation operator ^$\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. Subsequently, we need to learn a DA model $\mathcal{D}:\left(\mathbf{x}^{b}(t_{0}),\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\},\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}\right)\mapsto\mathbf{x}^{a}(t_{0})$, where the reanalysis serves as the reference.
@@ 4.2 Dataset
In this study, we utilized the DABench [33] dataset while incorporating AMSU-A and MHS satellite observations to train and evaluate the potential of XiChen for satellite assimilation tasks. DABench comprises 14 years of ERA5 reanalysis data, spanning from 2010 to 2023, as well as GDAS prepbufr observations, which were interpolated to a grid resolution of 1.40625 ($128\times 256$ grids). The observations are temporally interpolated to achieve a 3-hour temporal resolution interval, whereby measurements are systematically acquired at 0, 3, 6, 9, 12, 15, 18, and 21 hours UTC daily. Observations obtained within a 1.5-hour temporal window preceding and succeeding these designated temporal reference points are considered temporally coincident. We examined 5 atmospheric variables (each with 9 pressure levels) and four surface variables, resulting in a total of 49 variables. The atmospheric variables include geopotential (Z), temperature (T), specific humidity (Q), the zonal component of wind (U), and the meridional component of wind (V). The 9 sub-variables at different vertical levels are denoted by abbreviating their short names along with their respective pressure levels (e.g., Z500 indicates the geopotential at 500 hPa). The 4 surface variables consist of 2-meter temperature (T2M), 10-meter zonal component of wind (U10), 10-meter meridional component of wind (V10), and mean sea level pressure (MSLP). For further details on the dataset, we refer readers to the article [33].
For the AMSU-A and MHS observations, nearest neighbor interpolation was utilized to map both the raw satellite data and the auxiliary information onto a 1.40625 spatial resolution and 3-hour temporal resolution interval. In this study, we excluded observations from regions located above 60 north and south latitude to avoid complications due to sea ice  [48]. Additionally, we implemented a straightforward quality control procedure that eliminated data with brightness temperature values exceeding 350 K or falling below 150 K.
@@ 4.3 The fully AI-driven global weather forecasting system
The XiChen model is based on our proposed Conditional Hybrid Neural Operator (CHNO), which incorporates cross-attention to integrate conditional information into our Hybrid Neural Operator (HNO). This design enables a unified model architecture adaptable to multiple downstream tasks. Additionally, since the Adaptive Fourier Neural Operator (AFNO) [49] performs only global feature extraction through convolution operations in the Fourier space, we introduce a convolution-based branch to enhance local feature extraction capabilities, thereby developing the HNO. For further details on the CHNO and XiChen model architectures, refer to Figure S2. A comparison between HNO and AFNO in the context of medium-range weather forecasting is provided in the Supplementary Text and Figure S4.
@@
where $\mathbf{x}(t_{0})$ denotes the initial field to be optimized, $\mathbf{x}^{b}(t_{0})$ represents the background field, and $\mathbf{y}(t_{k})$ represents the observations at time $t_{k}$. The forecasting model, denoted as $\mathcal{M}_{t_{0}\rightarrow t_{k}}$, maps the initial field to the field at time $t_{k}$. Additionally, $\mathbf{B}$ and $\mathbf{R}$ represent the background and observation error covariance matrices, respectively. By combining the gradient $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ with the background field $\mathbf{x}^{b}(t_{0})$, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. Among them, our DA task involves sequentially employing DA models for different observational data. For example, in this study, XiChen first assimilates GDAS prepbufr data, then assimilates AMSU-A observations, and ultimately assimilates MHS observations. This configuration enhances XiChen’s scalability. The introduction of new observations in the future merely requires fine-tuning the respective observation operators and DA model components, eliminating the need to retrain the entire DA system. For details of cascade DA, please refer to the Supplementary Text and Figure S1 (b).
@@ 4.4 Pre-training for Medium-range Weather Forecasting
We aim to design a foundation model that can be pre-trained on the reanalysis data and then be finetuned to solve various downstream weather tasks to build a fully AI-driven global weather forecasting system. As illustrated in Figure 6(b), during the pre-training process, the foundation model takes the reanalysis field $\mathbf{x}(t)$ as the input and the lead time $\Delta t$ as the condition and outputs the predicted state $\hat{\mathbf{x}}(t+\Delta t)$. In this study, we sample $\Delta t$ from the set $\{1,3,6,12,24\}$ hours. This process enables the the foundation model to effectively capture the relationships among various weather variables while accounting for both the dynamical and thermodynamic characteristics of the atmosphere. To generate predictions for any time increments, the predictions from the learned forecasting model can be combined using the hierarchical temporal aggregation methodology proposed by Pangu-Weather [5]. Thus, the foundation model is pre-trained to minimize the following loss:
@@
where $\omega(v)$ denotes the pressure-weighted of variable $v$ [8], $\Delta\mathbf{x}(t+\Delta t)=\mathcal{M}_{t\rightarrow t+\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\Delta t)$ is the difference between the predicted state and the ground truth at time $t+\Delta t$, $\|\|_{1}$ represents the $L_{1}$-norm, and $L(i)$ is the latitude-weighting factor [36]:
@@
where $\text{lat}(i)$ is the latitude of the $i$th row of the grid. This term is commonly used in the training process of previous AI-based weather forecasting models [7, 5, 8]. This pre-training process costs about 45 hours to run on 8 A100-40GB GPUs.
@@
In practice, we first fine-tune the pre-trained forecasting model with roll-out steps from $T=2$ to $T=4$ to develop the XiChen-Short model. The weights obtained from the XiChen-Short model are then used to initialize the XiChen-Medium model, which is subsequently fine-tuned to achieve optimal forecasting performance for periods of 5 to 10 roll-out steps. We use the same sampled value of $\Delta t$ for all $T$ steps. Our forecasting model is similar to FuXi [10] in its forecasting process. Since the analysis field errors provided by the XiChen DA model are comparable to those of XiChen-Short’s 2-day forecasts initialized with ERA5, the 10-day medium-range weather forecasts using initial fields estimated by the XiChen DA model are generated by XiChen-Short for the first 3 days and by XiChen-Medium for the remaining 7 days. The fine-tuning process of the XiChen forecasting model costs about 94 hours to run on 8 A100-40GB GPUs. For additional information on the model training configuration, please refer to the Supplementary Text.

### v2

@@ 3 Discussion
The development of XiChen represents a significant breakthrough in fully AI-driven global weather forecasting. Leveraging the scalability of foundation models, XiChen is constructed by a pretrained medium-range forecasting model and fine-tuned to serve as observation operators and DA models. The entire pipeline of XiChen, from DA to 10-day weather forecasting, can be accomplished within only 15 seconds, representing an exceeding 400-fold acceleration compared to the operational NWP system [bauer2020ecmwf]. By integrating GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND observations, XiChen offers data-driven global weather forecasting capabilities and demonstrates superior performance compared to the GFS in both DA and forecasting tasks for several key variables. Notably, XiChen achieves a skillful lead time of over 8.75 days, which surpasses that of the GFS. Moreover, XiChen exhibits high scalability: when new satellite observations are introduced, it requires only fine-tuning the foundation model to the respective observation operator, followed by fine-tuning the corresponding assimilation model using the cascade DA framework proposed in this study. Similarly, if an observation becomes temporarily unavailable, the assimilation process can exclude the corresponding model from the pipeline. These capabilities are critical for advancing AI-based DA and forecasting toward practical operationalization. Finally, we verified that XiChen assimilation incorporates physical equilibrium constraints and flow-dependent characteristics through single-point perturbation DA experiments, further confirming its reliability. This achievement underscores the transformative potential of AI in advancing operational weather forecasting and highlights the necessity of sustained investments in AI research to tackle the complexities of Earth system modeling. As weather forecasting enters a new era, XiChen establishes a robust foundation for applying AI to deepen our forecasting of planetary systems, ultimately guiding decisions with profound implications for society and the environment. However, this does not imply that AI-driven global weather forecasting systems can fully replace NWP. Most AI-driven weather forecasting systems, including XiChen, rely heavily on reanalysis fields produced by NWP systems for training. Consequently, fully AI-driven global weather forecasting, exemplified by XiChen, serves as a valuable complement to NWP rather than a replacement.
@@
Although XiChen demonstrates significant potential for practical applications, it also exhibits certain limitations. First, the model experiences smoothing effects during both the DA and weather forecasting processes, a common challenge in AI-based weather forecasting models [bonavita2024some]. Addressing this issue requires further dedicated research. Potential solutions include designing more suitable loss functions [xu2024extremecast], incorporating physical constraints [xu2024weathergft], or employing post-processing techniques based on generative models [zhong2024fuxiextreme]. Second, the current version of XiChen has been trained and tested exclusively at a resolution of 1.40625. To predict finer-scale weather phenomena, future work should focus on extending this resolution to higher levels, such as 0.25 [ailabchen2025fengwu] or even 0.09 [han2024fengwughr]. Third, the current implementation of XiChen integrates satellite observations from only four satellite instruments—MHS, ASCAT, AMSU-A, and SATWND—significantly fewer than those utilized in operational NWP systems [zhong2024fuxiextreme]. In the future, incorporating additional indirect observational data could significantly enhance the performance of the analysis field. Finally, the model produces only deterministic weather forecasts. However, making critical decisions requires awareness of all possible weather scenarios and their associated probabilities. Therefore, it is necessary to extend the model to a probabilistic version. Addressing this limitation in the future may involve reconstructing the model to adopt a probabilistic framework [price2025probabilistic] or incorporating perturbations through the deterministic XiChen system [chen2024machine]. Ultimately, our goal is to develop a fully AI-driven global weather forecasting system that rivals the IFS HRES model in both assimilation and forecasting capabilities. Achieving this ambitious goal necessitates expertise in atmospheric science and the seamless integration of strengths from the atmospheric and AI communities to work together.
@@ 4.1 Notation and problem statement
We aim to develop a fully AI-driven global weather forecasting system that integrates both forecasting and DA models. For the forecasting model, given a system state $\mathbf{x}(t)$, our objective is to predict the state $\mathbf{x}(t+\Delta t)$ at a future time $t+\Delta t$. To accomplish this, the forecasting model $\mathcal{M}_{t\rightarrow t+\Delta t}:\mathbf{x}(t)\mapsto\hat{\mathbf{x}}(t+\Delta t)$ is trained to map the state of the atmosphere at present to a predicted state $\hat{\mathbf{x}}(t+\Delta t)$ for the future time $t+\Delta t$. Regarding the assimilation model, given a background field $\mathbf{x}^{b}(t_{0})$ at time $t_{0}$, GDAS prepbufr observations $\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\}$, and satellite observations $\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}$ during the DAW $t_{0},\cdots,t_{K}]$, our goal is to learn to estimate the optimal initial field $\mathbf{x}^{a}(t_{0})$ by fusing these data. To assimilate the satellite observations, we need to learn an observation operator $\hat{\mathcal{H}}^{s}:\left(\mathbf{x}(t),\theta^{s}(t)\right)\mapsto\mathbf{y}^{s}(t)$, where $\theta^{s}(t)$ denotes the auxiliary observational information at time $t$. Subsequently, we need to learn a DA model $\mathcal{D}:\left(\mathbf{x}^{b}(t_{0}),\{\mathbf{y}^{c}(t_{0}),\cdots,\mathbf{y}^{c}(t_{K})\},\{\mathbf{y}^{s}(t_{0}),\cdots,\mathbf{y}^{s}(t_{K})\}\right)\mapsto\mathbf{x}^{a}(t_{0})$, where the reanalysis serves as the reference.
@@ 4.2 Dataset
In this study, we utilized the DABench [wang2024benchmark] dataset while incorporating MHS, ASCAT, AMSU-A, and SATWND satellite observations to train and evaluate the potential of XiChen for satellite assimilation tasks. DABench comprises 14 years of ERA5 reanalysis data, spanning from 2010 to 2023, as well as GDAS prepbufr observations, which were interpolated to a grid resolution of 1.40625 ($128\times 256$ grids). The observations are temporally interpolated to achieve a 3-hour temporal resolution interval, whereby measurements are systematically acquired at 0, 3, 6, 9, 12, 15, 18, and 21 hours UTC daily. Observations obtained within a 1.5-hour temporal window preceding and succeeding these designated temporal reference points are considered temporally coincident. We examined 5 atmospheric variables (each with 9 pressure levels) and four surface variables, resulting in a total of 49 variables. The atmospheric variables include geopotential (Z), temperature (T), specific humidity (Q), the zonal component of wind (U), and the meridional component of wind (V). The 9 sub-variables at different vertical levels are denoted by abbreviating their short names along with their respective pressure levels (e.g., Z500 indicates the geopotential at 500 hPa). The 4 surface variables consist of 2-meter temperature (T2M), 10-meter zonal component of wind (U10), 10-meter meridional component of wind (V10), and mean sea level pressure (MSLP). For further details on the dataset, we refer readers to the article [wang2024benchmark].
For the AMSU-A and MHS observations, nearest neighbor interpolation was utilized to map both the raw satellite data and the auxiliary information onto a 1.40625 spatial resolution and 3-hour temporal resolution interval. In this study, we excluded observations from regions located above 60 north and south latitude to avoid complications due to sea ice  [noh2023assimilation]. Additionally, we implemented a straightforward quality control procedure that eliminated data with brightness temperature values exceeding 350 K or falling below 150 K. Since MHS observations are highly influenced by clouds, a basic cloud-detection method was applied to exclude cloud-contaminated portions of the MHS data. For details on the cloud-detection methodology, please refer to the reference paper [ma2013introduction]. For the 10-meter wind field data obtained from ASCAT and the upper-air wind field data obtained from SATWIND observations, we adopt an identical processing methodology to that implemented for GDAS prepbufr observational datasets.
@@
The TC ground truth used in this study is obtained from the IBTrACS dataset, accessible at here, which compiles recent and historical TC records from multiple agencies and provides the most comprehensive global archive of TC information.
@@ 4.3 The fully AI-driven global weather forecasting system
The XiChen model is based on our proposed Conditional Hybrid Neural Operator (CHNO), which incorporates cross-attention to integrate conditional information into our Hybrid Neural Operator (HNO). This design enables a unified model architecture adaptable to multiple downstream tasks. Additionally, since the Adaptive Fourier Neural Operator (AFNO) [guibas2021adaptive] performs only global feature extraction through convolution operations in the Fourier space, we introduce a convolution-based branch to enhance local feature extraction capabilities, thereby developing the HNO. For further details on the CHNO and XiChen model architectures, refer to Figure S2. A comparison between HNO and AFNO in the context of medium-range weather forecasting is provided in the Supplementary Text and Figure S4.
@@
where $\mathbf{x}(t_{0})$ denotes the initial field to be optimized, $\mathbf{x}^{b}(t_{0})$ represents the background field, and $\mathbf{y}(t_{k})$ represents the observations at time $t_{k}$. The forecasting model, denoted as $\mathcal{M}_{t_{0}\rightarrow t_{k}}$, maps the initial field to the field at time $t_{k}$. Additionally, $\mathbf{B}$ and $\mathbf{R}$ represent the background and observation error covariance matrices, respectively. By combining the gradient $\nabla_{\mathbf{x}^{b}(t_{0})}\mathcal{J}^{\text{4DVar}}(\mathbf{x}^{b}(t_{0}))$ with the background field $\mathbf{x}^{b}(t_{0})$, XiChen generates the analysis field. Utilizing the analysis field along with the lead time as inputs, XiChen produces either a forecast field or an updated background field. Among them, our DA task involves sequentially employing DA models for different observational data. For example, in this study, XiChen first assimilates GDAS prepbufr data, then assimilates MHS, ASCAT, AMSU-A, and SATWND observations. This configuration enhances XiChen’s scalability. The assimilation of new observations in the future merely requires fine-tuning the respective observation operators and DA model components, eliminating the need to retrain the entire DA system. For details of the cascade DA framework, please refer to the Supplementary Text and Figure S1 (b).
@@ 4.4 Pre-training for Medium-range Weather Forecasting
We aim to design a foundation model that can be pre-trained on the reanalysis data and then be finetuned to solve various downstream weather tasks to build a fully AI-driven global weather forecasting system. As illustrated in Figure 9(b), during the pre-training process, the foundation model takes the reanalysis field $\mathbf{x}(t)$ as the input and the lead time $\Delta t$ as the condition and outputs the predicted state $\hat{\mathbf{x}}(t+\Delta t)$. In this study, we sample $\Delta t$ from the set $\{1,3,6,12,24\}$ hours. This process enables the foundation model to effectively capture the relationships among various weather variables while accounting for both the dynamical and thermodynamic characteristics of the atmosphere. To generate predictions for any time increments, the predictions from the learned forecasting model can be combined using the hierarchical temporal aggregation methodology proposed by Pangu-Weather [bi2023accurate]. Thus, the foundation model is pre-trained to minimize the following loss:
@@
where $\omega(v)$ denotes the pressure-weighted of variable $v$ [lam2023learning], $\Delta\mathbf{x}(t+\Delta t)=\mathcal{M}_{t\rightarrow t+\Delta t}(\mathbf{x}(t))_{v,i,j}-\mathbf{x}_{v,i,j}(t+\Delta t)$ is the difference between the predicted state and the ground truth at time $t+\Delta t$, $\|\|_{1}$ represents the $L_{1}$-norm, and $L(i)$ is the latitude-weighting factor [rasp2020weatherbench]:
@@
where $\text{lat}(i)$ is the latitude of the $i$th row of the grid. This term is commonly used in the training process of previous AI-based weather forecasting models [kurth2023fourcastnet, bi2023accurate, lam2023learning]. This pre-training process costs about 45 hours to run on 8 A100-40GB GPUs.
@@
In practice, we first fine-tune the pre-trained forecasting model with roll-out steps from $T=2$ to $T=4$ to develop the XiChen-Short model. The weights obtained from the XiChen-Short model are then used to initialize the XiChen-Medium model, which is subsequently fine-tuned to achieve optimal forecasting performance for periods of 5 to 10 roll-out steps. We use the same sampled value of $\Delta t$ for all $T$ steps. Our forecasting model is similar to FuXi [fudanchen2023fuxi] in its forecasting process. Since the analysis field errors provided by the XiChen DA model are comparable to those of XiChen-Short’s 2-day forecasts initialized with ERA5, the 10-day medium-range weather forecasts using initial fields estimated by the XiChen DA model are generated by XiChen-Short for the first 3 days and by XiChen-Medium for the remaining 7 days. The fine-tuning process of the XiChen forecasting model costs about 94 hours to run on 8 A100-40GB GPUs. For additional information on the model training configuration, please refer to the Supplementary Text.